OpenStudy (anonymous):

can anyone help me? f(x)= (x-2)2 Find f(2)

hero (hero):

f(x) = (x-2)^2 To find f(2), everywhere you see x, replace it with 2, then evaluate: f(2) = (2-2)^2 f(2) = (0)^2 f(2) = 0

OpenStudy (anonymous):

I wrote the question wrong. f(x) = (x-2)^2 Find f(-2) Can you help me with that?

OpenStudy (anonymous):

OK f(x)=(x-2)^2 Now we are told to put the value -2 into the variable x and see what the function f(x) will give us when x=-2 . so: f(-2)=((-2)-2)^2 f(-2)=(-2-2)^2 f(-2)=(-4)^2 f(-2)=16

OpenStudy (anonymous):

I am not very good with math. english is my strong subject. I am trying to comprehend what I am doing. The answer that I got was -6. I do not understand how I got that. The way that my teacher explains it to me is so very confusing.

hero (hero):

It's good I always have fiddle around (wink wink) to trail me lol

OpenStudy (anonymous):

Another question? f(x) = x^4 + 2x^2 - 1 Find f(2)

OpenStudy (anonymous):

summer, first is this one - is it clear why for f(x)=(x-2)^2 f(-2)=16 ?

OpenStudy (anonymous):

No, sadly it is not.

OpenStudy (anonymous):

OK: f(-2)=(-2-2)^2 so far so good ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ok. how much is : (-2-2) ?

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

yes. that's right. nos how much is: (-4)^2 ?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

how did you get 6 before ?

OpenStudy (anonymous):

But u didnt do exponents first. I thought ur suppoose to do exponents first

OpenStudy (anonymous):

nope :) BEDMAS is an acronym that stands for; B-brackets E-exponents DM-multiply or divide (left to right) AS-add subtract (left to right) http://mathcentral.uregina.ca/QQ/database/QQ.09.07/h/brit1.html

OpenStudy (anonymous):

remember the word "BEDMAS" !!!! the B tells you that first you need to handle the Brackets.

OpenStudy (anonymous):

Ok what please excuse my dear aunt sally?

OpenStudy (anonymous):

same thing :) here the P stands for Paranthesis (another word for Brackets)

OpenStudy (anonymous):

Also here the Paranthesis are handled first, only then the Exponents.

OpenStudy (anonymous):

OK let me try another one of my problems. Do you have a few to wait and see if I get it correct?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

Thank you SOOOOOOO much :)

OpenStudy (anonymous):

f(x) =x^4 + 2x^2-1 f(2)

OpenStudy (anonymous):

OK - go ahead and start, step by step

OpenStudy (anonymous):

first - just rewrite it, and put "2" instead of every "x"

OpenStudy (anonymous):

even better - put "(2)" instead of every "x"

OpenStudy (anonymous):

done

OpenStudy (anonymous):

post what you did

OpenStudy (anonymous):

f(2) = (2)^4 + 2(2)^2-1

OpenStudy (anonymous):

OK, now we have parantheses here, but they just have the number 2 in them - so we don't have anything more we need to do with them. Next, we need to handle the exponents. Can you do that now ?

OpenStudy (anonymous):

16 and 4. Cuz (2)^4 = 16 and (2)^2 = 4

OpenStudy (anonymous):

correct - so now write the whole equation with those values ^^^

OpenStudy (anonymous):

16 + 4-1

OpenStudy (anonymous):

16 + (2)4 - 1?

OpenStudy (anonymous):

Almost, look: f(2) = (2)^4 + 2*(2)^2-1 f(2) = 16 + 2*4 -1

OpenStudy (anonymous):

yeah!!!

OpenStudy (anonymous):

OK, so we did paranthesis , then we did exponents. Next is multiplications and divisions.

OpenStudy (anonymous):

23?

OpenStudy (anonymous):

Write all the steps - it's a good drill to get it down ...

OpenStudy (anonymous):

Continue from here: f(2) = 16 + 2*4 -1

OpenStudy (anonymous):

so now we are doing multiplication (there is no division)

OpenStudy (anonymous):

16 + 8 -1= 23

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

congrats!

OpenStudy (anonymous):

Thanks! One more to make sure that I totally understand? I really really appreciate all the help!

OpenStudy (anonymous):

of course the last thing you did were the final Additions and Subtractions

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

YEAH!!

OpenStudy (anonymous):

that's the spirit!

OpenStudy (anonymous):

\[f(x)= 1/4x+2x Find f(8)\]

OpenStudy (anonymous):

Like this?